OpenStudy (anonymous):

The tides around Cherokee Bay range between a low of one foot to a high of five feet. The tide is at its lowest point when time, t, is 0 and completes a full cycle over a 24 hour period. What is the amplitude, period, and midline of a function that would model this periodic phenomenon?

OpenStudy (ddcamp):

Amplitude = 1/2 (Xmax - Xmin) Period = Time to complete a full cycle Midline = 1/2(Xmax + Xmin)